The balance sheets at the end of each of the first two years of operations indicate the following:

Kellman Company
Year 2 Year 1
Total current assets $611,431 $561,313
Total investments 63,987 51,189
Total property, plant, and equipment 866,804 702,453
Total current liabilities 119,450 85,844
Total long-term liabilities 284,513 228,043
Preferred 9% stock, $100 par 94,961 94,961
Common stock, $10 par 586,033 586,033
Paid-in capital in excess of par-common stock 68,057 68,057
Retained earnings 389,208 252,017

Using the balance sheets for Kellman Company, if net income is $109,787 and interest expense is $31,426 for Year 2, what is the return on stockholders' equity for Year 2 (round percent to two decimal points)?

a.11.26%

b.7.12%

c.9.88%

d.8.35%
Answer #1

Beginning stockholders' equity = Preferred 9% stock, $100 par + Common stock, $10 par + Paid-in capital in excess of par-common stock + Retained earnings

= 94,961 + 586,033 + 68,057 + 252,017

= $1,001,068

Ending stockholders' equity = Preferred 9% stock, $100 par + Common stock, $10 par + Paid-in capital in excess of par-common stock + Retained earnings

= 94,961 + 586,033 + 68,057 + 389,208

= $1,138,259

Average stockholders' equity = (Beginning stockholders' equity + Ending stockholders' equity)/2

= (1,001,068 + 1,138,259)/2

= $1,069,663.5

Return on stockholders' equity = Net income/Average stockholders' equity

= 109,787/1,069,663.5

= 10.26%

Correct answer is 10.26% but there is no option having answer as 10.26%.

Kindly comment if you need further assistance. Thanks